# RLVR-World: Training World Models with Reinforcement Learning

### v1

@@ Abstract
World models predict state transitions in response to actions and are increasingly developed across diverse modalities. However, standard training objectives such as maximum likelihood estimation (MLE) often misalign with task-specific goals of world models, i.e., transition prediction metrics like accuracy or perceptual quality. In this paper, we present RLVR-World, a unified framework that leverages reinforcement learning with verifiable rewards (RLVR) to directly optimize world models for such metrics. Despite formulating world modeling as autoregressive prediction of tokenized sequences, RLVR-World evaluates metrics of decoded predictions as verifiable rewards. We demonstrate substantial performance gains on both language- and video-based world models across domains, including text games, web navigation, and robot manipulation. Our work indicates that, beyond recent advances in reasoning language models, RLVR offers a promising post-training paradigm for enhancing the utility of generative models more broadly.
@@ 1 Introduction
World models [16, 30], which predict state transitions under action interventions, offer opportunities for autonomously evaluating and optimizing the behavior policies of intelligent agents [19, 20]. By scaling world models across various modalities–such as text [57, 8, 14], video [6, 1], and sensory data [46, 65]–significant progress has been made in a range of applications, including games [7, 27], robotics [64], and autonomous driving [22, 44].
Effectively scaling world models involves training highly expressive models, e.g., Transformers, on massive data. This critically depends on differentiable, rich, and stable training signals. In practice, world models are typically trained using surrogate objectives such as maximum likelihood estimation (MLE). For instance, language models are trained via next-token prediction, which supports reasoning through chain-of-thought generation prior to producing final answers [59], while diffusion models optimize a variational lower bound of the log-likelihood [21]. Moreover, non-end-to-end architectures that incorporate separately trained components like visual tokenizers [10, 43] enhance training efficiency and stability, especially for large-scale models. This paradigm has built powerful probabilistic world models, as well as many other foundation models [4].
However, the ultimate objectives of world models are beyond capturing complex distributions of data, but to meet the usage requirements of transition prediction metrics, such as high accuracy or perceptual quality. Surrogate and non-end-to-end optimizations are agnostic of and often diverge from this. In fact, training video models with standard mean squared error is known to produce blurry predictions [37], and in language models, likelihood-based objectives have been linked to issues like repetition and hallucination [31, 26, 56]. A promising emerging approach is tuning pre-trained models to directly optimize toward the target task via reinforcement learning with verifiable rewards (RLVR) [29, 15], which replaces the learned reward model in reinforcement learning from human feedback (RLHF) [40] with a faithful, rule-based reward function. Using this approach, the language model community is now producing impressive advances in tackling complex math reasoning and code generation problems.
@@
  • Language world models: Beyond the success of large language models (LLMs) in math and code domains, we introduce the world modeling task as a new testbed for RLVR in LLMs. This task, predicting the transition of verbal world states, naturally lends itself to using prediction accuracy as a verifiable reward. Our experiments show that RLVR can effectively fine-tune LLMs as language world models, yielding significant improvements, including $\bf+30.7\%$ accuracy on text-based game state prediction [57] and $\bf+15.1\%$ F1 score on web page state prediction [8].
  • Video world models: We pioneer the RLVR fine-tuning of autoregressive video world models [61] by directly measuring and optimizing perceptual metrics of decoded predicted frames against ground-truth observations. Notably, our method achieves substantial gains, e.g., $\bf+9.2\%$ relative improvement on LPIPS [68], on robot manipulation trajectory prediction [5], with merely a few hundred RLVR gradient steps, in contrast to the hundreds of thousands required by MLE training to achieve. We further prove the effectiveness of RLVR in bridging the gap between pre-trained models and the target world modeling task by showing that it mitigates the repetition issue, a phenomenon commonly seen in LLMs [31] and also observed in our pre-trained video world model.
Finally, we demonstrate the utility of reinforced world models in downstream applications, including policy evaluation [32] and model-predictive control [8]. We hope our method, experiments, and analysis will inspire future research to apply RLVR as a general post-training paradigm to significantly boost the usefulness of world models, and more broadly, generative models.
Figure 1: Training world models with reinforcement learning. (Left) World models are typically pre-trained using surrogate objectives like MLE, which misalign with task-specific prediction metrics. (Right) We propose post-training world models with RLVR to directly optimize these metrics.
@@ 2 Related Work
Learning accurate world models to predict environment state transitions in response to actions is fundamental to model-based planning and reinforcement learning [52]. Due to the inherent complexity and uncertainty of real-world environments, world models are more commonly implemented as generative models of next states conditioned on current states and actions [9, 16], rather than as deterministic models [45, 20]. For visual observations, sequential variational autoencoders have been widely adopted [25, 18, 17, 60], while recent advances have seen the emergence of diffusion-based world models [64, 2, 7], which offer superior visual fidelity. Another important line of work formulates world models autoregressively by discretizing visual inputs into sequences of tokens [39, 22, 61, 1]. This approach naturally extends to modeling additional modalities through unified token-based state representations. For example, previous work explores language-based state representation [57], which has been applied for building world models of the Internet in web agents [8, 14]. Similarly, trajectories of proprioceptive sensors and actuators can also be quantized into token sequences [46, 65]. Motivated by the prospect of building multimodal world models [28, 34] and the potential to leverage the large language model (LLM) ecosystem, our work explores the RLVR paradigm for training world models under the unified autoregressive formulation.
@@
Reinforcement learning has emerged as a critical paradigm for post-training generative models to better align with human preferences or task-specific objectives. In language models, InstructGPT [40] employs reinforcement learning from human feedback (RLHF) to enhance harmlessness, helpfulness, and honesty. However, RLHF is susceptible to reward model overoptimization [13]. In contrast, DeepSeek-R1 [48, 15] adopts reinforcement learning with verifiable rewards (RLVR), achieving significant advances in math, code, and logical reasoning domains. For visual generative models, text-to-image diffusion models have also been fine-tuned via reinforcement learning [3, 11] to optimize measurable metrics (e.g., compressibility) or human evaluations [62]. Our work identifies world modeling as an underexplored yet natural fit for RLVR, where prediction accuracy serves as a task-aligned, verifiable reward, enabling direct optimization of generative models across various modalities as world models.
@@ 3 Preliminaries
Given an image $x\in\mathbb{R}^{H\times W\times 3}$, the encoder of a discrete visual tokenizer [55] maps $x$ to its latent representation $h\in\mathbb{R}^{h\times w\times d}$. This latent is then quantized by performing a nearest neighbors lookup in a codebook of embeddings $C=\{e_{i}\}_{i=1}^{K}$, yielding a discrete representation $z\in[K]^{h\times w}$, which is passed through a decoder to reconstruct the original image $x$. The token map $z$ can be flattened into a 1D sequence of length $h\times w$ and subsequently modeled by autoregressive models such as decoder-only Transformers [10]. For videos in $\mathbb{R}^{T\times H\times W\times 3}$, a straightforward approach is to tokenize each frame independently using an image tokenizer [39, 7, 34]. However, this leads to excessively long token sequences. To mitigate this, Wu et al. [61] propose a compressive tokenization method that exploits temporal redundancy in videos by tokenizing each frame into a reduced number of $n$ tokens $z_{t}\in[K_{1}]^{n}$, conditioned on shared $N$ context tokens $z_{c}\in[K_{2}]^{N}$.
Group relative policy optimization (GRPO) [48]
is originally developed for post-training LLMs with reinforcement learning. Compared to PPO [47], GRPO eliminates the need for a value function and estimates advantages in a group-relative manner. Specifically, given a question $q$, GRPO samples a group of responses $\{o_{i}\}_{i=1}^{G}$ from the behavior policy $p_{\theta_{\text{old}}}$, and computes the advantage of each response by normalizing its reward $R_{i}$ within the group:
$\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G})}.$ (1)
@@
where $p_{\theta}^{i,t}$ denotes $p_{\theta}(o_{i,t}\mid q,o_{i,<t})$ for simpilicity. Refer to Shao et al. [48] for more details.
@@ 4.1 Problem Formulation
Environments simulated by world models are typically formulated as a Markov decision process (MDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},p,r,\gamma)$. Depending on the target task, the state space $\mathcal{S}$ can flexibly include various modalities, such as verbal, visual, or proprioceptive signals. At each timestep, the agent observes a state $s_{t}\in\mathcal{S}$, takes an action $a_{t}\in\mathcal{A}$, then transits to a new state according to the distribution $p(s_{t+1}\mid s_{t},a_{t})$, and receives an immediate reward $r_{t}\sim r(s_{t},a_{t})$. More generally, the states can be partially observable and the transitions can be modeled as a $k$-order Markov process $p(s_{t+1}\mid s_{t-k+1:t},a_{t-k+1:t})$  [^1] While partial observations are commonly denoted $o_{t}$ in classical RL literature, we keep using $s_{t}$ to avoid confusion with the generated outputs $\{o_{i}\}_{i=1}^{G}$ in our RLVR context.. World models need to learn to approximate the state transition $p$ and the reward function $r$ accurately. Since rewards can be considered as an extended dimension of the state space [9], our focus is on modeling the transition distribution $p(s_{t+1}\mid s_{t-k+1:t},a_{t-k+1:t})$.
@@ 4.2 World Models as Sequence Modeling
While different architectures have been proposed for world models on top of different modalities, next-token prediction by decoder-only Transformers has emerged as a general-purpose formulation applicable to tasks across various modalities. We unify world models into this general sequence modeling framework. To transform states and actions into tokens, different modalities have unique, commonly used tokenization schemes:
  • Languages are processed by standard text tokenization techniques like BPE [12].
  • Images and videos are encoded by learned visual tokenizers.
  • Low-dimensional continuous values, e.g., proprioceptive signals or joint torques, can be quantized to uniform bins over a fixed range.
@@ 4.3 Prediction Metrics as Verifiable Rewards
(1) Instead of solving the original environment MDP in Section 4.1, we focus on learning its transition distribution by formulating the next-state prediction process as another MDP, optimized using RLVR. (2) RLVR-World is a general framework, where input/output sequences and reward functions can be domain-specifically designed. We describe them in the experimental sections. (3) Our framework is compatible with various RL algorithms [47, 66], not limited to GRPO.
@@ 5.1 Text Game State Prediction
GPT-4 [57] 73.90% 51.60% 64.76%
Dataset and task.
We use ByteSized32-State-Prediction [57], a dataset of text game state transitions for evaluating how well LLMs can serve as text-based world simulators. The dataset contains 76,369 transitions from 31 distinct text games, with 2954 high-quality transitions selected for testing. In this task, an LLM models the world simulator function $F:\mathcal{C}\times\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}\times\mathcal{R}\times\mathcal{T}$, where $\mathcal{C}$ represents natural language contexts describing the task and action semantics, $\mathcal{S}$ is the state space encoded as JSON objects, $\mathcal{A}$ is the action space, $\mathcal{R}$ is the task reward, and $\mathcal{T}=\{0,1\}$ indicates task completion.
World model.
We use DeepSeek-R1-Distill-Qwen-1.5B [15, 63] as our base model. Due to its limited capability, we first apply supervised fine-tuning (SFT) using responses generated by DeepSeek-R1 [15], and then fine-tune with RLVR using either a binary accuracy reward $R=\mathbb{I}\left((\hat{s}^{\prime},\hat{r},\hat{w})=(s,r,w)\right)$, or a task-specific one to reflect the problem structure (see Appendix A.1).
Results.
The results are shown in Table 1. Since predicting state transitions is inherently more difficult when action changes the state, all models perform significantly better on unchanged cases. Compared to SFT, our RLVR-World with a minimalist binary reward can substantially improve performance, achieving +34.7% accuracy for unchanged and +8.9% for changed cases. Further incorporating human knowledge through a tailored reward yields even larger gains (+44.8% unchanged, +9.6% changed). These results enable our 1.5B model to rival the overall performance of GPT-4, although the accuracy on changed cases still falls short due to the limited base model’s capacity.

### v2

@@ Abstract
World models predict state transitions in response to actions and are increasingly developed across diverse modalities. However, standard training objectives such as maximum likelihood estimation (MLE) often misalign with task-specific goals of world models, i.e., transition prediction metrics like accuracy or perceptual quality. In this paper, we present RLVR-World, a unified framework that leverages reinforcement learning with verifiable rewards (RLVR) to directly optimize world models for such metrics. Despite formulating world modeling as autoregressive prediction of tokenized sequences, RLVR-World evaluates metrics of decoded predictions as verifiable rewards. We demonstrate substantial performance gains on both language- and video-based world models across domains, including text games, web navigation, and robot manipulation. Our work indicates that, beyond recent advances in reasoning language models, RLVR offers a promising post-training paradigm for enhancing the utility of generative models more broadly. Code, datasets, models, and video samples are available at the project website: https://thuml.github.io/RLVR-World.
Figure 1: Training world models with reinforcement learning. (Left) As world models adopt increasingly advanced and scalable architectures, they are typically pre-trained or supervised fine-tuned using surrogate objectives such as maximum likelihood estimation (MLE), which misalign with task-specific prediction metrics. (Right) We propose post-training world models via reinforcement learning with verifiable rewards (RLVR) to directly optimize for these metrics.
@@ 1 Introduction
World models ha2018recurrent ; lecun2022path , which predict state transitions under action interventions, offer opportunities for autonomously evaluating and optimizing the behavior policies of intelligent agents hafner2025mastering ; hansen2023td . By scaling world models across various modalities–such as text wang2024can ; chae2024web ; gu2024your , video videoworldsimulators2024 ; agarwal2025cosmos , and sensory data schubert2023generalist ; yin2025trajectory –significant progress has been made in a range of applications, including games bruce2024genie ; kanervisto2025world , robotics yang2023learning , and autonomous driving hu2023gaia ; russell2025gaia .
Effectively scaling world models involves training highly expressive models, e.g., Transformers, on massive data. This critically depends on differentiable, rich, and stable training signals. In practice, world models are typically trained using surrogate objectives such as maximum likelihood estimation (MLE). For instance, language models are trained via next-token prediction, which supports reasoning through chain-of-thought generation prior to producing final answers wei2022chain , while diffusion models optimize a variational lower bound of the log-likelihood ho2020denoising . Moreover, non-end-to-end architectures that incorporate separately trained components like visual tokenizers esser2021taming ; rombach2022high enhance training efficiency and stability, especially for large-scale models. This paradigm has built powerful probabilistic world models, as well as many other foundation models bommasani2021opportunities .
However, the ultimate objectives of world models are beyond capturing complex distributions of data, but to meet the usage requirements of transition prediction metrics, such as high accuracy or perceptual quality. Surrogate or non-end-to-end optimizations are often infeasible for, agnostic to, or even diverge from this. In fact, even with differentiable objectives beyond likelihood, non-end-to-end architectures like autoregressive models based on discrete tokenizers or diffusion models cannot directly optimize them. Typical likelihood objectives are not well aligned with the world modeling task: in language models, likelihood-based objectives have been linked to issues like repetition and hallucination holtzman2019curious ; li2023repetition ; kalai2024calibrated ; wang2020exposure , and training video models with standard mean squared error is known to produce blurry predictions mathieu2015deep . Moreover, the widely adopted training paradigm of teacher-forcing next-step prediction is also unaware of accumulation errors over multi-step horizons. A promising emerging approach is tuning pre-trained models to directly optimize toward the target task via reinforcement learning with verifiable rewards (RLVR) lambert2024t ; guo2025deepseek , which replaces the learned reward model in reinforcement learning from human feedback (RLHF) ouyang2022training with a faithful, rule-based reward function. Using this approach, the language model community is now producing impressive advances in tackling complex math reasoning and code generation problems.
@@
  • Language world models: Beyond the success of large language models (LLMs) in math and code domains, we introduce the world modeling task as a new testbed for RLVR in LLMs. This task, predicting the transition of verbal world states, naturally lends itself to using prediction accuracy as a verifiable reward. Our experiments show that RLVR can effectively fine-tune LLMs as language world models, yielding significant improvements, including $\bf+30.7\%$ accuracy on text-based game state prediction wang2024can and $\bf+15.1\%$ F1 score on web page state prediction chae2024web .
  • Video world models: We pioneer the RLVR fine-tuning of autoregressive video world models wu2024ivideogpt by directly measuring and optimizing perceptual metrics of decoded predicted frames against ground-truth observations. Notably, our method achieves substantial gains, e.g., $\bf+9.2\%$ relative improvement on LPIPS zhang2018unreasonable , on robot manipulation trajectory prediction brohan2022rt , with merely a few hundred RLVR gradient steps, in contrast to the hundreds of thousands required by MLE training to achieve. It also yields state-of-the-art performance compared with advanced world models zhou2024dino ; ko2023learning . We further prove the effectiveness of RLVR in bridging the gap between pre-trained models and the target world modeling task by showing that it mitigates the repetition issue, a phenomenon commonly seen in LLMs li2023repetition and also observed in our pre-trained video world model.
Finally, we demonstrate the utility of reinforced world models in downstream applications, including policy evaluation li2024evaluating and model-predictive control chae2024web . We hope our method, experiments, and analysis will inspire future research to apply RLVR as a general post-training paradigm to significantly boost the usefulness of world models, and more broadly, generative models.
@@ 2 Related Work
Learning accurate world models to predict environment state transitions in response to actions is fundamental to model-based planning and reinforcement learning sutton1998reinforcement . Due to the inherent complexity and uncertainty of real-world environments, world models are more commonly implemented as generative models of next states conditioned on current states and actions chua2018deep ; ha2018recurrent , rather than as deterministic models schrittwieser2020mastering ; hansen2023td . For visual observations, sequential variational autoencoders have been widely adopted kaiser2019model ; hafner2019learning ; hafner2019dream ; wu2023pre , while recent advances have seen the emergence of diffusion-based world models yang2023learning ; alonso2024diffusion ; bruce2024genie , which offer superior visual fidelity. Another important line of work formulates world models autoregressively by discretizing visual inputs into sequences of tokens micheli2022transformers ; hu2023gaia ; wu2024ivideogpt ; agarwal2025cosmos . This approach naturally extends to modeling additional modalities through unified token-based state representations. For example, previous work explores language-based state representation wang2024can , which has been applied for building world models of the Internet in web agents chae2024web ; gu2024your . Similarly, trajectories of proprioceptive sensors and actuators can also be quantized into token sequences schubert2023generalist ; yin2025trajectory . Motivated by the prospect of building multimodal world models kondratyuk2023videopoet ; liu2024world and the potential to leverage the large language model (LLM) ecosystem, our work explores the RLVR paradigm for training world models under the unified autoregressive formulation.
@@
Reinforcement learning has emerged as a critical paradigm for post-training generative models to better align with human preferences or task-specific objectives. In language models, InstructGPT ouyang2022training employs reinforcement learning from human feedback (RLHF) to enhance harmlessness, helpfulness, and honesty. However, RLHF is susceptible to reward model overoptimization gao2023scaling . In contrast, DeepSeek-R1 shao2024deepseekmath ; guo2025deepseek adopts reinforcement learning with verifiable rewards (RLVR), achieving significant advances in math, code, and logical reasoning domains. For visual generative models, text-to-image diffusion models have also been fine-tuned via reinforcement learning black2023training ; fan2023dpok to optimize measurable metrics (e.g., compressibility) or human evaluations xu2023imagereward . Our work identifies world modeling as an underexplored yet natural fit for RLVR, where prediction accuracy serves as a task-aligned, verifiable reward, enabling direct optimization of generative models across various modalities as world models.
@@ 3 Preliminaries
Given an image $x\in\mathbb{R}^{H\times W\times 3}$, the encoder of a discrete visual tokenizer van2017neural maps $x$ to its latent representation $h\in\mathbb{R}^{h\times w\times d}$. This latent is then quantized by performing a nearest neighbors lookup in a codebook of embeddings $C=\{e_{i}\}_{i=1}^{K}$, yielding a discrete representation $z\in[K]^{h\times w}$, which is passed through a decoder to reconstruct the original image $x$. The token map $z$ can be flattened into a 1D sequence of length $h\times w$ and subsequently modeled by autoregressive models such as decoder-only Transformers esser2021taming . For videos in $\mathbb{R}^{T\times H\times W\times 3}$, a straightforward approach is to tokenize each frame independently using an image tokenizer micheli2022transformers ; bruce2024genie ; liu2024world . However, this leads to excessively long token sequences. To mitigate this, Wu et al. wu2024ivideogpt propose a compressive tokenization method that exploits temporal redundancy in videos by tokenizing each frame into a reduced number of $n$ tokens $z_{t}\in[K_{1}]^{n}$, conditioned on shared $N$ context tokens $z_{c}\in[K_{2}]^{N}$.
Group relative policy optimization (GRPO) shao2024deepseekmath
is originally developed for post-training LLMs with reinforcement learning. Compared to PPO schulman2017proximal , GRPO eliminates the need for a value function and estimates advantages in a group-relative manner. Specifically, given a question $q$, GRPO samples a group of responses $\{o_{i}\}_{i=1}^{G}$ from the behavior policy $p_{\theta_{\text{old}}}$, and computes the advantage of each response by normalizing its reward $R_{i}$ within the group: $\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G})}.$
@@
where $p_{\theta}^{i,t}$ denotes $p_{\theta}(o_{i,t}\mid q,o_{i,<t})$ for simpilicity. Refer to Shao et al. shao2024deepseekmath for more details.
@@ 4.1 Problem Formulation
Environments simulated by world models are typically formulated as a Markov decision process (MDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},p,r,\gamma)$. Depending on the target task, the state space $\mathcal{S}$ can flexibly include various modalities, such as verbal, visual, or proprioceptive signals. At each timestep, the agent observes a state $s_{t}\in\mathcal{S}$, takes an action $a_{t}\in\mathcal{A}$, then transits to a new state according to the distribution $p(s_{t+1}\mid s_{t},a_{t})$, and receives an immediate reward $r_{t}\sim r(s_{t},a_{t})$. More generally, the states can be partially observable and the transitions can be modeled as a $k$-order Markov process $p(s_{t+1}\mid s_{t-k+1:t},a_{t-k+1:t})$  [^1] While partial observations are commonly denoted $o_{t}$ in classical RL literature, we keep using $s_{t}$ to avoid confusion with the generated outputs $\{o_{i}\}_{i=1}^{G}$ in our RLVR context.. World models need to learn to approximate the state transition $p$ and the reward function $r$ accurately. Since rewards can be considered as an extended dimension of the state space chua2018deep , our focus is on modeling the transition distribution $p(s_{t+1}\mid s_{t-k+1:t},a_{t-k+1:t})$.
@@ 4.2 World Models as Sequence Modeling
While different architectures have been proposed for world models on top of different modalities, next-token prediction by decoder-only Transformers has emerged as a general-purpose formulation applicable to tasks across various modalities. We unify world models into this general sequence modeling framework. To transform states and actions into tokens, different modalities have unique, commonly used tokenization schemes: languages are processed by standard text tokenization techniques like BPE gage1994new ; images and videos are encoded by learned visual tokenizers; and low-dimensional continuous values, e.g., proprioceptive signals, can be quantized to uniform bins over a fixed range.
@@ 4.3 Prediction Metrics as Verifiable Rewards
(1) Instead of solving the original environment MDP in Section 4.1, we focus on learning its transition distribution by formulating the next-state prediction process as another MDP, optimized using RLVR. (2) RLVR-World is a general framework, where input/output sequences and reward functions can be domain-specifically designed. We describe them in the experimental sections. (3) Our framework is compatible with various RL algorithms schulman2017proximal ; yu2025dapo , not limited to GRPO.
@@ 5.1 Text Game State Prediction
Dataset and task.
We use ByteSized32-State-Prediction wang2024can , a dataset of text game state transitions for evaluating how well LLMs can serve as text-based world simulators. The dataset contains 76,369 transitions from 31 distinct text games, with 2954 high-quality transitions selected for testing. In this task, an LLM models the world simulator function $F:\mathcal{C}\times\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}\times\mathcal{R}\times\mathcal{T}$, where $\mathcal{C}$ represents natural language contexts describing the task and action semantics, $\mathcal{S}$ is the state space encoded as JSON objects, $\mathcal{A}$ is the action space, $\mathcal{R}$ is the task reward, and $\mathcal{T}=\{0,1\}$ indicates task completion.
World model.
We use DeepSeek-R1-Distill-Qwen-1.5B and 7B guo2025deepseek ; yang2024qwen25 as our base model. Due to their limited capability, we first apply supervised fine-tuning (SFT) using responses generated by DeepSeek-R1 guo2025deepseek , and then fine-tune with RLVR using either a binary accuracy reward $R=\mathbb{I}\left((\hat{s}^{\prime},\hat{r},\hat{w})=(s,r,w)\right)$, or a task-specific one to reflect the problem structure (see Appendix A.1).
Results.
The results are shown in Table 1. Since predicting state transitions is inherently more difficult when action changes the state, all models perform significantly better on unchanged cases. For the 1.5B base model, our RLVR-World with a minimalist binary reward can substantially improve performance over SFT, achieving +34.7% accuracy for unchanged and +8.9% for changed cases. Incorporating human knowledge through a tailored reward yields even larger gains (+44.8% for unchanged, +9.6% for changed). When scaled to the 7B base model, our method achieves overall performance surpassing that of GPT-4, although the accuracy on changed cases still falls short due to the limited base model’s capacity.
@@ Results.
Base
(R1-Distill-Qwen-7B)
46.90% 5.53% 29.92%
SFT 65.94% 31.32% 51.76%
RLVR-World
(Ours, binary)
83.08% 40.33% 65.53%
GPT-4 wang2024can 73.90% 51.60% 64.76%